OpenStudy (anonymous):

Sketch a graph of y=f(x) f(x)=1/4x+9

OpenStudy (anonymous):

I just need someone to help walk me through this!

OpenStudy (anonymous):

Is that\[\frac{ x }{ 4 }+9\]or\[\frac{ 1 }{ 4x }+9\]Can't tell from the way you wrote it.

OpenStudy (anonymous):

The second one, sorry!

OpenStudy (anonymous):

there is a screen shot of the problem

OpenStudy (anonymous):

np. Notice that you will have a vertical asymptote at x=0 and as x approaches 0 from the right, f(x) will go to +infinity.. as x approaches 0 from the left, f(x) will go to -infinity

OpenStudy (anonymous):

I'll have to go off of your question, since I can't open up the document since I don't have Word. Also, you will have a y-intercept at (-1/36, 0)

OpenStudy (anonymous):

I can send it via paint if you have that

OpenStudy (anonymous):

Just so you understand what I am doing.

OpenStudy (anonymous):

I have paint. jpegs work, too. You also have a horizontal asymptote at y = 9 = f(x) because as x goes to either +-infinity, the f(x) goes to 9.

OpenStudy (anonymous):

so 0,9

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What is 0,9 ?

OpenStudy (anonymous):

Your graph will look roughly like this.